# How to set AdamW’s weight decay as you scale model and dataset size

### v2

@@ 1 Introduction
Perhaps more importantly, the EMA perspective tells us how to set the weight decay parameter. Specifically, we can take into account the dataset and batch sizes to measure the timescale in epochs, i.e. $\tau_{\text{epoch}}=\tau_{\text{iter}}/M$, where $M=N/\mathrm{B}$ denotes the number of iterations per epoch in a training dataset of $N$ samples under a batch size of $\mathrm{B}$, which measures the fraction of the dataset that contributes to the current weights through the EMA. We empirically observed that the optimal $\tau_{\text{epoch}}$ is stable across model and dataset sizes, which, consequently provides a guideline for scaling weight decay. In particular, as the dataset size $N$ increases, we find that the optimal weight decay falls (Sec. 4.1). Moreover, if we follow the usual µP recommendation to decrease the learning rate $\eta$ as the model width increases, we find that the optimal weight decay should increase as the model size increases (Sec. 4.2). We validate both of these predictions in ResNet, vision transformer trained on CIFAR-10 and ImageNet, and NanoGPT pre-training on OpenWebText.
@@
  • The optimal weight decay increases with model width.
@@ 2 Background
we know that $y_{i}\sim\sqrt{\text{fan\_in}}\times W^{0}_{ij}$, so to ensure that $y_{i}\sim 1$, we need $W^{0}_{ij}\sim 1/\sqrt{\text{fan\_in}}$. However, this square root scaling only arises in very specific circumstances, when each term in the sum, $W^{0}_{ij}x_{j}$, is zero-mean and uncorrelated. These requirements hold for the initial weights, as they are sampled IID from a zero-mean distribution. However, these conditions do not hold for the update (the precise reason why is complex; see Yang et al., 2022, for details),
@@ 3.1 AdamW as an EMA
Now we will show that these weight updates (Eq. 4) can be understood as an EMA [^2] For an introduction to EMAs, see en.wikipedia.org/wiki/Exponential_smoothing. Recall that a generic exponential moving average estimate, $\text{ema}_{t}$ can be written as,
@@ 4 Scaling weight decay across model and dataset sizes
For all experiments, we implemented the model using PyTorch and performed optimization using its AdamW implementation. For all tasks, we did not use weight decay on the normalization layers. Experiments were conducted on an internal cluster of Nvidia TitanX/1080ti/2080ti/L40s GPUs, and only one GPU card was used at a time for each run. The results for image classification tasks are the averages from three distinct random seeds, for LLM pre-training tasks we only use one seed due to resource constraints. Comprehensive details on the hyperparameter search range and model specifications can be found in Appendix F.
@@ 4.2 Transferring weight decay across model sizes
Figure 4: For LLM pre-training, optimal weight decay increases with model size whereas the optimal timescales transfer. We trained 8-layer GPTs on OpenWebText, with various widths (line colors), under two $\eta_{\text{base}}$, with actual learning rates for each width scaled following µP with $s=\tfrac{1024}{\textrm{width}}$. We plot the final validation loss against various $\lambda_{\text{base}}$ (dots on lines, varied by powers of $2$). We align all lines to the rightmost point for better visibility, where the numbers in the bracket denote the actual optimal validation loss at each width. We considered training under two parameterizations, if we keep $\lambda$ decoupled from $s$ then the optimal $\lambda_{\text{base}}$ shows a clear shift with model size (top rows), if $\lambda$ increases with $s$, optimal $\lambda_{\text{base}}$ becomes much stable across widths.
@@
Importantly, scaling the weight decay correctly has important implications even for the original µP predictions about how the optimal learning rate transfers across model sizes. Specifically, we trained a ResNet-18 on CIFAR-10 and on 320K images downscaled from ImageNet, with a fixed $\lambda_{\text{base}}=1.0$ and swept $\eta_{\text{base}}$ from $10^{-5}$ to $10^{-2}$. In the top row of Fig. 5A and B, we use the standard µP scaling (Eq. 13), with a fixed weight decay. Remarkably, we found that the optimal base learning rate varied dramatically across model sizes. This indicates that µP scaling of the optimal learning rate breaks down in AdamW. Indeed, the original µP paper (Yang et al., 2022) did not examine AdamW, and this finding is confirmed by recent work on hyperparameter transfer in large-scale transformers (Lingle, 2024; Blake et al., 2024). We hypothesized that this breakdown was due to the timescale changing with model size (Eq. 14), and thus we could restore the usual µP scaling of the learning rate by increasing the weight decay with model size (Eq. 15) to fix the timescale when sweeping $\eta_{\text{base}}$. To confirm this finding, in the bottom row of Fig. 5A and B, we increase the weight decay with model size (Eq. 15). We found that the optimal base learning rate was now far more stable when varying model sizes, confirming our hypothesis. We observed similar behavior when training ViTs of different widths on the ImageNet subset, the results are presented in Appendix D, and follow-up work, citing us, observed similar behavior in large-scale training of transformers (Blake et al., 2024).
@@ 5 Related work
There have been a number of follow-up works that follow up on the concepts presented in this paper and cite us. First, Blake et al. (2024) confirmed that increasing weight decay as we increase model width fixes learning rate transfer. Second, Anonymous (2024) has built on our EMA interpretation of AdamW to understand and propose better learning rate schedules. Third, Bjorck et al. (2024) looked at the scaling of optimal learning rates with dataset size. They found that the optimal learning rate decreases with dataset size ($N$), with a fixed batch size, and a fixed weight decay. This broadly what you would expect from our findings, that you need to increase the EMA timescale as the dataset size increases, and you can do that by reducing the weight decay, or reducing the learning rate. However, they found that the dependence of learning rate on dataset size was sublinear. We suspect this is because as you modify the learning rate with weight decay fixed, you simultaneously change two things: the EMA timescale (which you probably do want to change) and the behaviour of the optimizer close to initialization, which you probably do not want to change (Appendix A).
@@
Second, our viewpoint (Sec. 3.2) suggests that the effects of the learning rate, $\eta$, on the magnitude of the updates relative to the random initialization should become less relevant than the EMA timescale as training proceeds (Appendix A). While this seems strange (of course the learning rate matters alot), there are some suggestions in the literature that indeed the EMA timescale may be more important in some settings. Specifically, Wortsman et al. (2024) ran proxies for large-scale LLM pretraining. Fig. 6 in their paper showed that if $\gamma=\eta\lambda=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ is fixed, the final validation loss is relatively insensitive to the learning rate, $\eta$. In contrast, if $\lambda$ is fixed, the validation loss shows much more pronounced sensitivity to the learning rate, $\eta$. This makes sense if the key hyperparameter is $\tau_{\text{iter}}$, as modifying the learning rate, $\eta$, with $\lambda$ fixed implies $\tau_{\text{iter}}=1/(\eta\lambda)$ also changes. Indeed, the $\gamma=\eta\lambda$ parameterization is often used in SGD, where it is identified as the effective learning rate (Zhang et al., 2019b; Li & Arora, 2019; Li et al., 2020; Wan et al., 2021; Li et al., 2022). Recent works extend the effective learning rate perspectives to AdamW by viewing AdamW as sign gradient descent (D’Angelo et al., 2024) or as rotation speed on a fixed radius sphere (Kosson et al., 2023). However, these literautres do not study how optimal $\eta\lambda$ transfers across model and dataset sizes.
@@
Third, Lingle (2024) assessed the accuracy of µP in large-scale LLM pretraining. Remarkably, they found that the usual µP scaling of the optimal learning rate with model size broke down for AdamW. We provide an explanation and fix for this by noting that µP theory considers only the behaviour of the first few optimization steps, relative to the random initialization. The issue is that if the weight decay, $\lambda$, is fixed, then $\tau_{\text{iter}}$ changes with model size (as $\tau_{\text{iter}}$ depends on $\eta$, and $\eta$ changes with model size following the usual µP recommendations). We propose a fix for issue by proposing to strengthen weight decay as the model size increases, and validate the fix experimentally.
@@
Fourth, many papers have considered scaling with batch size (which we hold fixed) (Zhang et al., 2019a; Malladi et al., 2022; Wang & Aitchison, 2024). These papers find that when the batch size is small, the optimal learning rate is also small. Then, as the batch size increases, the optimal learning rate also increases. However, there’s a limit, the critical batch size, at which point increases in the batch size no longer allow you to increase the learning rate. Once you hit the critical batch size, you no longer get any benefit from increasing the batch size, and you instead must increase the number of iterations (McCandlish et al., 2018; Shallue et al., 2019). Note that while very recent work suggests that the optimal batch size does increase with dataset size, they confirm that this increase is sublinear, implying that the number of training iterations will increase (Zhang et al., 2024). As soon as you need to increase the number of iterations, our recommendation to hold $\tau_{\text{epoch}}$ constant comes into play.
@@ A.3.3 Decoupling the learning rates for the batch/layernorm parameters
It turns out that we already do something different with these parameters. In particular, it is common practice333 We cannot find literature explicitly studying the effect of dropping weight decay for normalization layers’ parameters, however, its importance has been noticed in vision tasks, e.g. by Jia et al. (2018) and https://github.com/JiahuiYu/slimmable_networks/issues/15, as well as in language model training: https://github.com/karpathy/minGPT/blob/master/mingpt/model.py#L227 to optimize them using the same learning rate, $\eta$, as the other parameters, but to drop weight decay. This standard setting is depicted in Fig. 7E and Fig. 8E, and we can see that $\eta$ does still change the behavior of the optimizer.
@@ Appendix B The weights in an EMA
Appendix C Weight norm

### v3

@@ 1 Introduction
Perhaps more importantly, the EMA perspective tells us how to set the weight decay parameter. Specifically, we can take into account the dataset and batch sizes to measure the timescale in epochs, i.e. $\tau_{\text{epoch}}=\tau_{\text{iter}}/M$, where $M=N/\mathrm{B}$ denotes the number of iterations per epoch in a training dataset of $N$ samples under a batch size of $\mathrm{B}$, which measures the fraction of the dataset that contributes to the current weights through the EMA. We empirically observed that the optimal $\tau_{\text{epoch}}$ is stable across model and dataset sizes, which, consequently provides a guideline for scaling weight decay. In particular, under a fixed learning rate schedule, as the dataset size $N$ increases, we find that the optimal weight decay falls (Sec. 4.1). Moreover, if we follow the usual µP recommendation to decrease the learning rate $\eta$ as the model width increases, we find that the optimal weight decay should increase as the model size increases (Sec. 4.2). We validate both of these predictions in ResNet, vision transformer trained on CIFAR-10 and ImageNet, and NanoGPT pre-training on OpenWebText.
@@
  • The optimal weight decay increases with model width under µP scaling of learning rate.
@@ 2 Background
we know that $y_{i}\sim\sqrt{\text{fan\_in}}\times W^{0}_{ij}$, so to ensure that $y_{i}\sim 1$, we need $W^{0}_{ij}\sim 1/\sqrt{\text{fan\_in}}$. However, this square root scaling only arises in very specific circumstances, when each term in the sum, $W^{0}_{ij}x_{j}$, is zero-mean and uncorrelated. These requirements hold for the initial weights, as they are sampled I.I.D from a zero-mean distribution. However, these conditions do not hold for the update (the precise reason why is complex; see Yang et al., 2022, for details),
@@ 3.1 AdamW as an EMA
Now we will show that these weight updates (Eq. 4) can be approximately understood as an EMA [^2] For an introduction to EMAs, see en.wikipedia.org/wiki/Exponential_smoothing. Recall that a generic exponential moving average estimate, $\text{ema}_{t}$ can be written as,
@@
Notice that AdamW’s optimization process differs slightly from a standard EMA process. Standard EMA typically assumes the updates $q_{t}$’s are independent, while the $q_{t}$’s (Eq. 6) in AdamW’s EMA view are correlated since later updates depend on the weight $\operatorname{ema}_{t}$, which is affected by the current update $q_{t}$. However, we can still use EMA as an approximation, and it provides useful insight and intuition for understanding the hyperparameters.
@@ 4 Scaling weight decay across model and dataset sizes
For all experiments, we implemented the model using PyTorch and performed optimization using its AdamW implementation. For all tasks, we did not use weight decay on the normalization layers. Experiments were conducted on an internal cluster of Nvidia TitanX/1080ti/2080ti/L40s GPUs, and only one GPU card was used at a time for each run. The results for image classification tasks are the averages from three distinct random seeds, for LLM pre-training tasks we only use one seed due to resource constraints. Comprehensive details on the hyperparameter search range and model specifications can be found in Appendix G.
@@ 4.1 Transferring weight decay across dataset sizes
Note that in the experiments considered, we used a fixed initial learning rate under different dataset sizes. Some recent works (Bjorck et al., 2024; Schaipp et al., 2025) suggest that, under a fixed weight decay and batch size, the optimal learning rate decreases as training run gets longer. We empirically verified this observation in our experiment setting (top rows in Fig. 12 AB in appendix). However, we also notice that if we scale $\lambda_{\text{base}}\propto 1/N$, the optimal $\eta$ becomes consistent across dataset sizes (bottom rows in Fig. 12 AB in appendix), indicating that scaling $\eta$ may not be necessary if we keep $\tau_{\text{epoch}}$ constant through adjusting $\lambda$ when $N$ scales.
@@ 4.2 Transferring weight decay across model sizes
Figure 4: For LLM pre-training, optimal weight decay increases with model size whereas the optimal timescales transfer. We trained 8-layer GPTs on OpenWebText, with various widths (line colors), under two $\eta_{\text{base}}$, with actual learning rates for each width scaled following µP with $s=\tfrac{1024}{\textrm{width}}$. We plot the final validation loss against various $\lambda_{\text{base}}$ (dots on lines, varied by powers of $2$). We align all lines to the rightmost point for better visibility, where the numbers in the brackets denote the actual optimal validation loss at each width. We considered training under two parameterizations, if we keep $\lambda$ decoupled from $s$ then the optimal $\lambda_{\text{base}}$ shows a clear shift with model size (top rows), if $\lambda$ increases with $s$, optimal $\lambda_{\text{base}}$ becomes much stable across widths.
@@
Importantly, scaling the weight decay correctly has important implications even for the original µP predictions about how the optimal learning rate transfers across model sizes. Specifically, we trained a ResNet-18 on CIFAR-10 and on 320K images downscaled from ImageNet, with a fixed $\lambda_{\text{base}}=1.0$ and swept $\eta_{\text{base}}$ from $10^{-5}$ to $10^{-2}$. In the top row of Fig. 5A and B, we use the standard µP scaling (Eq. 13), with a fixed weight decay. Remarkably, we found that the optimal base learning rate varied dramatically across model sizes. This indicates that µP scaling of the optimal learning rate breaks down in AdamW. Indeed, the original µP paper (Yang et al., 2022) did not examine AdamW, and this finding is confirmed by recent work on hyperparameter transfer in large-scale transformers (Lingle, 2024; Blake et al., 2024). We hypothesized that this breakdown was due to the timescale changing with model size (Eq. 14), and thus we could restore the usual µP scaling of the learning rate by increasing the weight decay with model size (Eq. 15) to fix the timescale when sweeping $\eta_{\text{base}}$. To confirm this finding, in the bottom row of Fig. 5A and B, we increase the weight decay with model size (Eq. 15). We found that the optimal base learning rate was now far more stable when varying model sizes, confirming our hypothesis. We observed similar behavior when training ViTs of different widths on the ImageNet subset, the results are presented in Appendix E, and follow-up work, citing us, observed similar behavior in large-scale training of transformers (Blake et al., 2024).
@@ 5 Related work
5.1 Work after us
There have been a number of follow-up works that follow up on the concepts presented in this paper. First, Blake et al. (2024) confirmed that increasing weight decay as we increase model width fixes learning rate transfer. Second, Bergsma et al. (2025b) has built on our EMA interpretation of AdamW to understand and propose better learning rate schedules. Third, Bjorck et al. (2024) looked at the scaling of optimal learning rates with dataset size. They found that the optimal learning rate decreases with dataset size ($N$), with a fixed batch size, and a fixed weight decay. This broadly what you would expect from our findings, that you need to increase the EMA timescale as the dataset size increases, and you can do that by reducing the weight decay, or reducing the learning rate. However, they found that the dependence of learning rate on dataset size was sublinear. We suspect this is because as you modify the learning rate with weight decay fixed, you simultaneously change two things: the EMA timescale (which you probably do want to change) and the behaviour of the optimizer close to initialization, which you probably do not want to change (Appendix A). Additionally, our experiments verified that $\tau_{\text{epoch}}$ transfers across model widths on small-scale problems while recent works  (Dey et al., 2025; Narayan et al., 2025) further conclude that $\tau_{\text{epoch}}$ is also transferrable in billion-parameter LLM pre-training.
Lastly, very recent work by Bergsma et al. (2025a) considered two critical problems unstudied in our current manuscript: The effect of batch size as well as the optimal value and transferability of $\tau_{\text{epoch}}$ when you only train the model for one epoch, as in LLM pre-training settings. In our experiments, we kept the batch size $B$ fixed when tuning other hyperparameter, however in real world pre-training, batch size $B$ is often tuned extensively to ensure maximum utilization of parallelism, Therefore it is important to understand the how to adjust other hyperparameter as we scale $B$. Fortunately, the heuristic of keeping $\tau_{\text{epoch}}$ (Eq. (10)) constant provides a guidance on how to achieve this: In order to keep $\tau_{\text{epoch}}$ unchanged as $B$ scales, one can have $\eta\propto B$ as suggested by McCandlish et al. (2018), however the picture is often more complicated in practice due to e.g. AdamW’s effective learning rate induced by the gradient normalizer (Malladi et al., 2022), as well as the potential risk of optimization diverging due to learning rate set too large. However, there exists an alternative approach to keep $\tau_{\text{epoch}}$ unchanged as we scale $B$, that is to have $\lambda\propto B$. Indeed Bergsma et al. (2025a) demonstrates the feasibility of such an approach, which shows optimal $\lambda$ scales linearly with respect to $B$, and allows for much more stability and flexibility compared with adjusting $\eta$ when scaling $B$. Additionally, our manuscript only considered high epoch training where the same dataset is repeated multiple times, and suggests that optimal $\tau_{\text{epoch}}$ should stay constant and lies between one and the total number of epochs as dataset sizes increase. However, this may not be the case, LLM pre-training’s one epoch setting, where each training data point is only presented once. Particularly, Bergsma et al. (2025a) show that optimal $\tau_{\text{epoch}}$ should not stay constant in one epoch setting, but rather follows a power law $\tau_{\text{epoch}}^{\textrm{optimal}}\approx(\rm TPP)^{-0.527}$, where the optimal $\tau_{\text{epoch}}$ decreases as the token-per-parameter (TPP) increases, and is always smaller than $1$ when $\rm TPP>1$. Intuitively, this suggests that when the training run is short, it would be better to average over all past updates (high $\tau_{\text{epoch}}$), whereas as training gets longer, it becomes more optimal to forget noisy updates seen in the beginning of the training (low $\tau_{\text{epoch}}$).
5.2 Work before us
@@
Second, our viewpoint (Sec. 3.2) suggests that the effects of the learning rate, $\eta$, on the magnitude of the updates relative to the random initialization should become less relevant than the EMA timescale as training proceeds (Appendix A). While this seems strange (of course the learning rate matters a lot), there are some suggestions in the literature that indeed the EMA timescale may be more important in some settings. Specifically, Wortsman et al. (2024) ran proxies for large-scale LLM pretraining. Fig. 6 in their paper showed that if $\gamma=\eta\lambda=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ is fixed, the final validation loss is relatively insensitive to the learning rate, $\eta$. In contrast, if $\lambda$ is fixed, the validation loss shows much more pronounced sensitivity to the learning rate, $\eta$. This makes sense if the key hyperparameter is $\tau_{\text{iter}}$, as modifying the learning rate, $\eta$, with $\lambda$ fixed implies $\tau_{\text{iter}}=1/(\eta\lambda)$ also changes. However, different from our work, Wortsman et al. (2024) do not explicitly study how $\tau_{\text{iter}}$ transfers across model sizes or if keeping $\lambda$ fixed would break µP. Indeed, the $\gamma=\eta\lambda$ parameterization is often used in SGD, where it is identified as the effective learning rate (Zhang et al., 2019b; Li & Arora, 2019; Li et al., 2020; Wan et al., 2021; Li et al., 2022). Recent works extend the effective learning rate perspectives to AdamW by viewing AdamW as sign gradient descent (D’Angelo et al., 2024) or as rotation speed on a fixed radius sphere (Kosson et al., 2023) [^3] As pointed out by one of the anonymous reviewers, the timescale parameter in an EMA process can been seen as a measurement of the relative update size studied by Kosson et al. (2023), see Appendix. C for detailed discussion.. However, these literatures do not study how optimal $\eta\lambda$ transfers across model and dataset sizes.
@@
Third, Lingle (2024) assessed the accuracy of µP in large-scale LLM pretraining. Remarkably, they found that the usual µP scaling of the optimal learning rate with model size broke down for AdamW. We provide an explanation and fix for this by noting that µP theory considers only the behaviour of the first few optimization steps, relative to the random initialization. The issue is that if the weight decay, $\lambda$, is fixed, then $\tau_{\text{iter}}$ changes with model size (as $\tau_{\text{iter}}$ depends on $\eta$, and $\eta$ changes with model size following the usual µP recommendations). We propose a fix for the issue by proposing to strengthen weight decay as the model size increases, and validate the fix experimentally.
@@
Fourth, many papers have considered scaling with batch size (which we hold fixed) (Zhang et al., 2019a; Malladi et al., 2022; Wang & Aitchison, 2024). These papers find that when the batch size is small, the optimal learning rate is also small. Then, as the batch size increases, the optimal learning rate also increases. However, there’s a limit, the critical batch size, at which point increases in the batch size no longer allow you to increase the learning rate. Once you hit the critical batch size, you no longer get any benefit from increasing the batch size, and you instead must increase the number of iterations (McCandlish et al., 2018; Shallue et al., 2019). Note that while very recent work suggests that the optimal batch size does increase with dataset size, they confirm that this increase is sublinear, implying that the number of training iterations will increase (Zhang et al., 2025). As soon as you need to increase the number of iterations, our recommendation to hold $\tau_{\text{epoch}}$ constant comes into play.
@@ 6 Conclusions
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.3.3 Decoupling the learning rates for the batch/layernorm parameters
It turns out that we already do something different with these parameters. In particular, it is common practice444 We cannot find literature explicitly studying the effect of dropping weight decay for normalization layers’ parameters, however, its importance has been noticed in vision tasks, e.g. by Jia et al. (2018) and https://github.com/JiahuiYu/slimmable_networks/issues/15, as well as in language model training: https://github.com/karpathy/minGPT/blob/master/mingpt/model.py#L227 to optimize them using the same learning rate, $\eta$, as the other parameters, but to drop weight decay. This standard setting is depicted in Fig. 7E and Fig. 8E, and we can see that $\eta$ does still change the behavior of the optimizer.
@@ Appendix B The weights in an EMA
Appendix C Connection between EMA timescale and relative update size
Consider an EMA given by $a_{t+1}=(1-\gamma)a_{t}+\gamma b_{t}$. Assume that the $b_{t}$s are independently random and identically distributed over time with mean zero, and focus on the relative updates in the steady state (large $t$). The relative update size for the scalar case is defined as
$r=\sqrt{\frac{\mathbb{E}[(a_{t+1}-a_{t})^{2}]}{\mathbb{E}[a_{t}^{2}]}}.$ (43)
To compute $r$, we need to determine $\mathbb{E}[a_{t}^{2}]$. We start by expressing $a_{t}$ in terms of $b_{t}$ using the EMA formula:
$a_{t}=\gamma\sum_{k=0}^{t-1}(1-\gamma)^{k}b_{t-1-k}+(1-\gamma)^{t}a_{0}.$ (44)
For simplicity, we assume $a_{0}=0$ in the steady state, leading to:
$a_{t}=\gamma\sum_{k=0}^{t-1}(1-\gamma)^{k}b_{t-1-k}.$ (45)
Squaring both sides gives:
$a_{t}^{2}=\gamma^{2}\left(\sum_{k=0}^{t-1}(1-\gamma)^{k}b_{t-1-k}\right)^{2}.$ (46)
Expanding the square, we have:
$a_{t}^{2}=\gamma^{2}\left(\sum_{k=0}^{t-1}(1-\gamma)^{2k}b_{t-1-k}^{2}+2\sum_{0\leq k<j<t}(1-\gamma)^{k+j}b_{t-1-k}b_{t-1-j}\right).$ (47)
Due to the independence and zero mean assumption of $b_{t}$, the cross terms $b_{t-1-k}b_{t-1-j}$ for $k\neq j$ have zero expectation. Thus, we have:
$\mathbb{E}[a_{t}^{2}]=\gamma^{2}\sum_{k=0}^{t-1}(1-\gamma)^{2k}\mathbb{E}[b_{t-1-k}^{2}].$ (48)
Assuming $\mathbb{E}[b_{t}^{2}]=\sigma^{2}$ and using the geometric series sum formula for large $t$, we get:
$\mathbb{E}[a_{t}^{2}]=\frac{\gamma^{2}\sigma^{2}}{1-(1-\gamma)^{2}}.$ (49)
This simplifies to:
$\mathbb{E}[a_{t}^{2}]=\frac{\gamma^{2}\sigma^{2}}{\gamma(2-\gamma)}.$ (50)
Next we will simplify $\mathbb{E}[(a_{t+1}-a_{t})^{2}]$, we begin by expanding the inner term
$(a_{t+1}-a_{t})^{2}=\gamma^{2}(b_{t}^{2}-2b_{t}a_{t}+a_{t}^{2}).$ (51)
Taking expectations, we have:
$\mathbb{E}[(a_{t+1}-a_{t})^{2}]=\gamma^{2}(\mathbb{E}[b_{t}^{2}]-2\mathbb{E}[b_{t}a_{t}]+\mathbb{E}[a_{t}^{2}]).$ (52)
Assuming $b_{t}$ and $a_{t}$ are independent and $b_{t}$ has mean zero, $\mathbb{E}[b_{t}a_{t}]=0$. Thus:
$\mathbb{E}[(a_{t+1}-a_{t})^{2}]=\gamma^{2}(\mathbb{E}[b_{t}^{2}]+\mathbb{E}[a_{t}^{2}]).$ (53)
Using $\mathbb{E}[b_{t}^{2}]=\sigma^{2}$ and substituting $\mathbb{E}[a_{t}^{2}]$ from the previous derivation:
$\mathbb{E}[(a_{t+1}-a_{t})^{2}]=\gamma^{2}\left(\sigma^{2}+\frac{\gamma^{2}\sigma^{2}}{\gamma(2-\gamma)}\right).$ (54)
This simplifies to:
$\mathbb{E}[(a_{t+1}-a_{t})^{2}]=\gamma^{2}\sigma^{2}\left(1+\frac{\gamma}{2-\gamma}\right).$ (55)
Finally, simplifying further gives:
$\mathbb{E}[(a_{t+1}-a_{t})^{2}]=\frac{2\gamma^{2}\sigma^{2}}{2-\gamma}.$ (56)
Thus, the expression for $r$ becomes:
$r=\sqrt{\frac{2\gamma\sigma^{2}}{\sigma^{2}}}=\sqrt{2\gamma},$ (57)
which indicates that the timescale parameter in the EMA is directly linked with the relative update size.
Appendix D Weight norm